OpenStudy (anonymous):

traffic signs are regulated by the manual on uniform traffic control devices (mutcd). the perimeter of some traffic sign must be 104 inches. also its length is 12inches longer than its width. find the dimensions of this sign

OpenStudy (stormfire1):

P = 104 = 2L + 2W L = 12 + W 104 = 2(12+W) + 2W 104 = 24 + 2W + 2W Solving for W you get: 80 = 4W W = 20 Since L = 12 + W, L = 32